OpenStudy (anonymous):

two men painted a room together in 3 hours. How long will it take each one working alone if the faster one can do the job in 2 1/2 hours less than the other one?

OpenStudy (anonymous):

thanks in advance for those who will answer :))

OpenStudy (jamesj):

The way in general to approach these problems is to think in terms of rates of doing something. Man 1 can paint r1 of a room in one hour; Man 2 paints at the rate r2. This means that Man 1 by himself can paint a room in 1/r1 hours and Man 2 in 1/r2 hours. So together they paint a room at a rate of r1 + r2. This means that the time required to paint one room is 1/(r1 + r2) From the problem we know that it takes 3 hours for them to paint a room together. hence 1/(r1 + r2) = 3. We also know that 1/r1 - 1/r2 = 2.5 [we have assumed Man 1 paints slower than Man 2] So now you have two equations in terms of r1 and r2 and it's a question of algebra to get them back in linear form so you can solve them. Can you take it from here?

OpenStudy (anonymous):

let A and B arae 2 men let A can finish in x hours so B can finish in x-5/2 hours in 1 hour x can do 1/x part of work in 1 hour B can do 1/(x+2.5) part of work togather in one hour they can do 1/x + 1/(x+2.5) part of work they finished togather in 3 hours so 3[1/x + 1/(x+2.5)] = 1 (3/x)+(3/(x+2.5))=1 3(x+2.5)+3(x)=(x)(x+2.5) 3x+7.5+3x=x^2+2.5x x^2-3.5x-7.5=0 2x^2-7x-15=0 (2x+3)(x-5)=0 x=-3/2 and x=5 time cannot be negative so x=5 x+2.5=7.5

OpenStudy (anonymous):

hope it was helpful!